OpenStudy (sstradic):

A paper popcorn cone measures 4 inches in diameter and is 9 inches high. What is the volume of popcorn that can be contained within this cone? Use 3.14 for pi. Round your answer to the nearest hundredth. 37.68 cubic inches 39.45 cubic inches 40.23 cubic inches 41.90 cubic inches

OpenStudy (mathstudent55):

\(\Large V_{cone} = \dfrac{1}{3}\pi r^2 h\)

OpenStudy (mathstudent55):

Use the formula above to find the volume.

OpenStudy (sstradic):

can u explain the formula pls

OpenStudy (mathstudent55):

\(\pi\) is 3.14 as your problem tells you. r is the radius of the cone. You have a diameter of 4 in., so your radius is half of that, or 2 in. h is the height of the cone which is 9 in. Use the numbers above in the formula and evaluate the expression.